The circuit shown has three identical resistors in series with device for measuring current (an ammeter). The value of the emf shown is 9.99 V,9.99 V, and each resistor has a value of 120 Ω.120 Ω.

Part A: What is the total resistance of this circuit (?eq)?(Req)?

?eq=Req=

ΩΩ

Part B: What is the value of the current (?)(I) flowing out of the emf?

?=I=

A

Part C: What is the value of the voltage drop (Δ?)(ΔV) across the second resistor?

Δ?=ΔV=
Answer #1

Resistance of each resistor, R = 120 Ω

Emf of the battery, V = 9.99 V

Part A

Total resistance when three identical resistors are connected in series,

Req = R + R + R = 3 x 120 = 360 Ω

Part B

Current flowing out of the battery, I = V/R = 9.99/360 = 0.028 A

Part C

Voltage drop across the second resistor, ΔV = IR = 0.028 x 120 = 3.36 V